# Event2Vec: Processing neuromorphic events directly by representations in vector space

### v1

@@ Abstract
The neuromorphic event cameras have overwhelming advantages in temporal resolution, power efficiency, and dynamic range compared to traditional cameras. However, the event cameras output asynchronous, sparse, and irregular events, which are not compatible with mainstream computer vision and deep learning methods. Various methods have been proposed to solve this issue but at the cost of long preprocessing procedures, losing temporal resolutions, or being incompatible with massively parallel computation. Inspired by the great success of the word to vector, we summarize the similarities between words and events, then propose the first event to vector (event2vec) representation. We validate event2vec on classifying the ASL-DVS dataset, showing impressive parameter efficiency, accuracy, and speed than previous graph/image/voxel-based representations. Beyond task performance, the most attractive advantage of event2vec is that it aligns events to the domain of natural language processing, showing the promising prospect of integrating events into large language and multimodal models. Our codes, models, and training logs are available at https://github.com/fangwei123456/event2vec.
@@ 2 View of Encoding
    Every element is a combination of an index and a position. Each word has a unique index in the dictionary. Converting words to indices is implemented by the tokenizer in NLP tasks, and the indices in Fig.1 are generated by the LLama-3 [23] tokenizer. The position of a word is determined by the location in the sentence, e.g., 00 for ”I” in the sentence ”I am glad to see you”. The index of an event is the tuple $(x_{i},y_{i},p_{i})$. However, the position is not $i$ but $t_{i}$, which indicates the temporal location in the event stream.
@@
3 Represent Event in Vector Spase
3.1 Event to Vector
@@
Given the similarities between words and events, it is natural to represent events in vector space, i.e., event to vector (event2vec). For an event $(x,y,t,p)$ obtained from an event camera with $H\times W$ spatial resolution, we embed it to the unique embedding 1D-index $i$ and embedding vector $\textbf{v}\in\mathbb{R}^{d}$:
$i$ $=p\cdot H\cdot W+y\cdot W+x,$ (1)
v $=\text{Embed}(i),$ (2)
where $\text{Embed}(\cdot)$ is the embedding layer with learnable weights $\textbf{W}\in\mathbb{R}^{(2\cdot H\cdot W)\times d}$, and $d$ is the embedding dimension.
A key difference between words and events is their spatial structure. The word is embedded in a 1D flattened sentence, while the event occurs in a 2D plane. However, the naive event2vec, Eq.(1)-(2) does not consider this feature. Thus, we modify the naive event2vec to the 2D event2vec formulation:
$i$ $=p\cdot H+y,$ (3)
$j$ $=p\cdot W+x,$ (4)
v $=[\text{Embed}_{y}(i),\text{Embed}_{x}(j)].$ (5)
Eq.(3) and Eq.(4) first encode $y$ and $x$ to unique indices, then Eq.(5) concatenates the embedding tensors.
3.2 Neighbor Semantics
It is worth noting that the real physical world is continuous, but the pixel coordinates captured by a camera are discrete. The generation of $(x,y)$ for an event introduces noise and quantization. Say, when capturing an identical object twice, we may get $(x,y)$ the first time and $(x+1,y)$, $(x,y-1)$ or other neighbor coordinates the second time. Thus, we expect that events with close coordinates should have similar semantics. Compared to naive event2vec Eq.(1)-(2) which can only output the same v when both $x$ and $y$ are the same for two events, the 2D event2vec Eq.(3)-(5) will output the same $\textbf{v}_{y}$ or $\textbf{v}_{x}$ when input events have either same $y$ or $x$. However, it still does not fully account for the similarity between neighbors, such as $(x,y)$ and $(x-1,y-1)$. Although neural networks have the potential to learn this spatial-semantic similarity without any inductive bias, learning may be difficult, i.e., may require many training epochs and large datasets.
To achieve neighbor similarity, a trivial idea is downsampling the spatial resolution. For example, if two events have the coordinates $(0,0)$ and $(1,1)$ at $128\times 128$ resolution, their coordinates will be both $(0,0)$ after downsampling the resolution to $64\times 64$. In fact, this method is similar to patchifying in ViT which uses grids rather than pixels as the element.
The second solution is involving the neighbors of an event manually. When processing $(x,y)$, we regard it as a group of coordinates and embed both the central point $(x,y)$ and its neighbors. Denote the number of neighbors as $n$, we expand the 2D event2vec as:
$\textbf{v}_{y}$ $=\sum_{k=-n}^{k=n}w_{k}\cdot\text{Embed}_{y}(p\cdot H+y+k),$ (6)
$\textbf{v}_{x}$ $=\sum_{k=-n}^{k=n}w_{k}\cdot\text{Embed}_{x}(p\cdot W+x+k),$ (7)
v $=[\textbf{v}_{y},\textbf{v}_{x}],$ (8)
where $w_{k}$ is a learnable parameter and initialized by the distance:
$w_{k}=\exp\Big{(}\frac{-k^{2}}{2}\Big{)}.$ (9)
With this neighbor embedding, the semantics of an event is determined not only by itself but also by its neighbors. This embedding is more consistent with the 2D spatial structure of the events.
3.3 Temporal Positional Encoding
Compared with words, events have explicit positions, i.e. timestamps. Transformers use sinusoidal or rotational positional encoding methods to encode position information[22, 24]. To incorporate temporal information from events, we can also use temporal positional encoding. Positional encoding in NLP tasks can be regarded as encoding the positions of words $\{0,1,...,L-1\}$, where $L$ is the sequence length. For events, the only thing to do is replace the inputs by $\{t_{0},t_{1},...,t_{L-1}\}$. Additionally, timestamps can also be added to embedding vectors directly because they already have explicit physical semantics, i.e. the physical time of events occurring.
@@ 4 Experiments
We validate the proposed event2vec methods on the neuromorphic ASL-DVS [16] classification task. The ASL-DVS dataset contains 24 letters (classes) from the American Sign Language. Each class contains 4200 samples and each sample lasts about 100 $ms$. We use the first 80% of the dataset as the train set and the last 20% as the test set. The train-test ratio is same as [16]. To avoid any challenges with code reproducibility and allow fair comparison for researchers using our methodology, we do not split randomly.
4.1 Training Hyparameters
We use a vanilla transformer encoder structure with only two layers, 256 features, 16-head self-attention, and 512 features in the feedforward layers. Note that the space complexity of self-attention is $\mathcal{O}(L^{2})$, where $L$ is the sequence length. Thus, we cannot process event streams with too many events. We just randomly sample $L=255$ events from each sample. The sampling is implemented by the choice function of Numpy [8], and the randomness is controlled by the random generator in Numpy. Although the random sampling will cause undetermined test accuracy, we find that the results are stable after about averaging over four repeated tests, which is also the test accuracy reported in this article. We set all random seeds as 0. We use the AdamW [25] optimizer with a learning rate of 0.001, batch size = 128, and cosine annealing learning rate scheduler [26]. We train the model for 64 epochs. We downsample the spatial size from $180\times 240$ to $32\times 48$.
We directly add the timestamps to the embedding vectors and do not use any additional temporal positional encoding techniques. We find that this simple encoding method works well enough. We normalize the timestamps because their values are too large and may cause numerical instability in neural networks. The maximum value of the timestamp in the train set is $t_{max}=521217$. For both train and test sets, we normalize the timestamp by $t=\frac{t}{t_{max}}$. We regard that the absolute values of timestamps are meaningless, and we set all timestamps in an event stream minus the first timestamp.
The experiments are carried out to show the effectiveness of our proposed event2vec, rather than seeking state-of-the-art accuracy. Thus, we do not use any data augmentation or regularization such as dropout and weight decay.
4.2 Accuracy, Efficiency, and Speed
With $n=2$ neighbors and 2D event2vec embedding, our model achieves 99.68% test accuracy. Due to random sampling, the model only utilizes about 0.00143% of all events in the training set in one iteration. After 64 epochs (40319 iterations), only 58% of the train set is used to train the model. The high test accuracy shows the extreme learning efficiency of the model equipped with our event2vec representation. Tab.1 compares the event2vec representation with some previous representations on the ASD-DVS dataset. It is worth noting that our method achieves high accuracy with an extremely small model size and does not require any event data preprocessing. Our model achieves 2185 samples/s training speed and 4320 samples/s inference speed on a Red Hat Enterprise Linux 8 server with Intel(R) Xeon(R) Gold 6240 CPU @ 2.60GHz, a single Nvidia Tesla V100-SXM2-32GB GPU, and 64GB RAM. For comparison, [27] takes 16.7 $ms$ to process one sample, and the execution time does not include the event to frame and voxel-graph preprocessing. Our method only requires 1000/4320=0.23 $ms$ for one sample, which is about 72$\times$ faster than [27]. Note that our speed includes the data reading and sampling.
Fig.2 compares the test accuracy of different event2vec methods. Note that we transform the training iteration steps to the ratio of used training data for ease of reading. The results show that the spatial structure of the 2D event2vec is vital, without which the model can hardly learn the relationship between close coordinates, leading to much lower accuracy. With the weighted neighbor embedding, the accuracy of 2D event2vec is further promoted, but the accuracy saturates with more than 2 neighbors. Thus, using $n=2$ neighbors is sufficient.
Fig.3 shows the test accuracy under different downsampling sizes and event2vec methods. When using the naive event2vec, lower spatial resolution is helpful for clustering neighbor events and leads to higher accuracy. When using the 2D event2vec with weighted neighbors, the neighbor semantics are inherent, and the quantization of coordinates under low spatial resolution may cause accuracy loss, e.g., accuracy under $9\times 12$ resolution is slightly lower than that under $36\times 48$ when using 2D event2vec, but higher when using the naive event2vec.
@@ 4.2 Accuracy, Efficiency, and Speed
Method Preprocessing Accuracy (%) Model Size (MB) Epoch
GNN+CNN [16] To graphs 90.1 19.46 150
GNN + Transformer [27] To images and voxel-graph 99.6 220.3 150
Event2vec + Transformer (Ours) None 99.68 4.13 64
Table 1: Comparison with previous representations on the ASL-DVS dataset.
(a) Effectiveness of 2D embedding
(b) Effectiveness of neighbors embedding
Figure 2: Effectiveness event2vec methods. (a) The spatial structure of the 2D event2vec is vital. (b) The induction of neighbors is helpful, but the effectiveness is marginal. In general, $n=2$ neighbors are good enough.
(a) Naive event2vec
(b) 2D event2vec with 2 neighbors.
Figure 3: Comparison of downsampling sizes. (a) Downsampling is helpful for the naive event2vec. (b) But its effectiveness is marginal, even harmful for 2D event2vec with neighbors.

### v2

@@ Abstract
The neuromorphic event cameras have overwhelming advantages in temporal resolution, power efficiency, and dynamic range compared to traditional cameras. However, the event cameras output asynchronous, sparse, and irregular events, which are not compatible with mainstream computer vision and deep learning methods. Various methods have been proposed to solve this issue, but at the cost of long preprocessing procedures, losing temporal resolutions, or being incompatible with massively parallel computation. Inspired by the great success of the word to vector, we summarize the similarities between words and events, then propose the first event to vector (event2vec) representation. We validate event2vec on classifying the DVS Gesture and ASL-DVS dataset, showing impressive learning efficiency and accuracy. Beyond task performance, the most attractive advantage of event2vec is that it aligns events to the domain of natural language processing, showing the promising prospect of integrating events into large language and multimodal models. Our codes, models, and training logs are available at https://github.com/fangwei123456/event2vec.
@@ 2 View of Encoding
    Every element is a combination of an index and a position. Each word has a unique index in the dictionary. Converting words to indices is implemented by the tokenizer in NLP tasks, and the indices in Fig.1 are generated by the LLama-3 [23] tokenizer. The position of a word is determined by the location in the sentence, e.g., 0 for ”I” in the sentence ”I am glad to see you”. The index of an event is the tuple $(x_{i},y_{i},p_{i})$. However, the position is not $i$ but $t_{i}$, which indicates the temporal location in the event stream.
@@
3 Represent Event in Vector Space
@@
Given the similarities between words and events, it is natural to represent events in vector space, i.e., event to vector (event2vec). For an event $(x,y,t,p)$ obtained from an event camera with $H\times W$ spatial resolution, we regard $(x,y,p)$ as the spatial coordinate and $t$ as the temporal position. We embed them to the embedding vector $\textbf{v}\in\mathbb{R}^{d}$:
v $=\text{Embed}_{s}(x,y,p)+\text{Embed}_{t}(t),$ (1)
where $\text{Embed}_{s}(\cdot)$ is the spatial embedding module and $\text{Embed}_{t}(\cdot)$ is the temporal encoding module. Eq.(1) fuses the spatial and temporal information by addition, which is also the most widely used fashion in LLMs.
3.1 Spatial Embedding
A trivial formulation for the spatial embedding module is using the standard embedding layer for word2vec, which is implemented efficiently by a look-up-table:
$\text{Embed}_{s}(x,y,p)=\textbf{W}[p\cdot H\cdot W+y\cdot W+x],$ (2)
where $\textbf{W}\in\mathbb{R}^{(2\cdot H\cdot W)\times D}$ is the learnable embedding matrix and $D$ is the embedding dimension. Eq.(2) maps each coordinate to a unique row index and selects the corresponding row from the embedding matrix.
The standard embedding layer does not preset any inductive bias between indices, and expects the model to learn relations between indices automatically. This design philosophy works well for words, for which the order of indices does not provide any information. For instance, we can not expect that the $i$-th word in a dictionary has a similar semantics to the $(i+1)$-th word. However, it is not the case for event coordinates.
Since an image is a continuous two-dimensional function [24], it is widely known that pixels in a small neighborhood are similar. Thus, we expect that events with close coordinates should have similar semantics:
$\text{Embed}_{s}(x+\Delta x,y+\Delta y,p+\Delta p)-\text{Embed}_{s}(x,y,p)\approx\textbf{0},$ (3)
when $\Delta x,\Delta y,\Delta p]\approx\textbf{0}$.
However, Eq.(2) does not consider the neighbor semantics, which may cause difficulty for learning. To solve this issue, possible solutions include pre-training the embedding layer to learn neighborhood semantics, adding a neighborhood loss term, or using a weighted sum of neighbors centered by $(x,y,p)$ as $\text{Embed}_{s}(x,y,p)$. Unfortunately, the former two methods require extra training procedures and parameter fine-tuning. The latter method causes computational costs to increase with the number of neighbors.
To achieve neighborhood semantics, we propose an elegant parametric algorithm 1 to generate the embedding matrix W. As the algorithm 1 shows, we first generate an arrangement c, then convert it to the probe tensor $\textbf{x},\textbf{y},\textbf{p}]\in\mathbb{N}^{(2\cdot H\cdot W)\times 3}$, which enumerates all event coordinates. Finally, we send the probe tensor to the parametric neural network $\phi$, and get the output embedding matrix $\textbf{W}\in\mathbb{R}^{(2\cdot H\cdot W)\times D}$.
Algorithm 1 The parametric algorithm for generating the embedding matrix
1:The height $H$, and width $W$ of events; the parametric neural network $\phi$
[…]
When using W generated from the algorithm 1 in Eq.(2), for an arbitrary event coordinate $(x,y,p)$, it can be found that:
$\textbf{W}[p\cdot H\cdot W+y\cdot W+x]=\phi(x,y,p).$ (4)
It is worth noting that $\phi$ is continuous and differentiable. According to the Taylor series, we have:
$\phi(x+\Delta x,y+\Delta y,p+\Delta p)-\phi(x,y,p)=$
$[\Delta x,\Delta y,\Delta p]\cdot[\frac{\partial\phi}{\partial x}(x,y,p),\frac{\partial\phi}{\partial y}(x,y,p),\frac{\partial\phi}{\partial p}(x,y,p)]^{\rm{T}}+o^{n},$ (5)
where $o^{n}$ represents the remainder term. Eq.(5) shows that when $\Delta x,\Delta y,\Delta p]\approx\textbf{0}$, then
$\phi(x+\Delta x,y+\Delta y,p+\Delta p)-\phi(x,y,p)\approx\textbf{0}+o^{n}\approx\textbf{0},$ (6)
indicating that the embedding tensors of $(x+\Delta x,y+\Delta y,p+\Delta p)$ and $(x,y,p)$ are similar. Thus, we incorporate neighborhood semantics into the spatial embedding module by the parametric neural network $\phi$.
3.2 Temporal Encoding
Timestamps represent the occurrence time of events, whose role is similar to the indices of words in a sentence. Recently, relative positional encoding methods [25, 26] have attracted increasing interest and gradually substitute the absolute positional encoding method, such as the sinusoidal encoding [22] and the positional-wise learnable encoding [27].
However, applying relative positional encoding methods to timestamps is not reasonable. These encoding methods are designed for discrete and uniform indices, while timestamps are nearly continuous and uneven. To solve this issue, we employ a neural network to learn temporal encoding from the difference of timestamps.
More specifically, the temporal encoding is implemented by stacked convolutional layers, taking the one-order time difference t[i+1]t[i]t[i+1]-t[i] for the $i$-th event in an event stream as inputs. Such a design brings the following advantages:
  1. (1)
    It utilizes the relative temporal distances of events directly, enabling the time-shift invariance.
  2. (2)
    It is a variant of position-wise learnable encoding but avoids the length generalization problem by taking continuous values as inputs.
  3. (3)
    The convolution operations involve adjacent events to generate the temporal encoding value of the center event, which is consistent with the neighborhood semantics.
@@ 4 Experiments
We validate the proposed event2vec methods on the neuromorphic DVS Gesture [28] and ASL-DVS [16] classification tasks. The DVS Gesture dataset contains 11 hand gesture categories and provides an official training/test split with 1176/288 samples. The ASL-DVS dataset contains 24 letters (classes) from the American Sign Language. Each class contains 4200 samples, and each sample lasts about 100 $ms$. We use the first 80% of the dataset as the training set and the last 20% as the test set. The train-test ratio is the same as [16]. To avoid any challenges with code reproducibility and allow fair comparison for researchers using our methodology, we do not split randomly.
4.1 Network Structure
We employ a multi-stage network structure to classify events, which is shown in Fig.2. The network contains the event2vec module, the linear attention backbone, and a linear layer as the final classification head.
4.1.1 Event Embedding
As shown in Fig.2, the spatial embedding module $\phi$ contains stacks of linear layers. It gradually increases the number of features from 3 to $D$. The layer normalization [29] layers are also inserted to stabilize training. Note that $\phi$ generates the embedding matrix W from the probe tensor as Algorithm 1, and then event coordinates are converted to tensors by Eq.(2). We can also use $\phi$ to process event coordinates and get the same embedding tensors directly, as shown by Eq.(4). We can choose the indirect or the direct method flexibly according to their training speed.
The temporal encoding module has a similar structure to the spatial embedding module, except that it uses depth-wise convolutional layers with a kernel size of 3 and a stride of 1.
Figure 2: The network structure for classifying event datasets with event2vec.
4.1.2 Backbone
Eq.(1) embeds an event stream to a sequence, which can be processed by RNNs, Transformers, or temporal convolutional networks. The number of events in one event stream is also the length of tokens $L$, which is much longer than that of words in NLP tasks. For example, the shortest sample in the DVS Gesture training dataset contains 35,267 events.
The space complexity of self-attention in Transformers is $\mathcal{O}(L^{2})$, where $L$ is the sequence length. Thus, it is impractical to process a long event stream with the vanilla Transformers. Fortunately, linear attentions [30, 31, 32] with the $\mathcal{O}(L)$ space complexity have been proposed to solve the quadratic complexity problem. In this article, we employ the gated linear attention [33] to process event embedding tensors.
It is worth noting that linear attentions can also be regarded as RNNs [34]. To enhance the long-term learning ability of the gated linear attention for processing extremely long event streams, we extend it to a bi-directional formulation. Denote $\textbf{S}[t]$ as the hidden states at the time-step $t$, the recurrence of $\textbf{S}[t]$ in linear attentions obeys the following rule:
$\textbf{S}[t]=f(\textbf{S}[t-1],\textbf{k}[t],\textbf{v}[t]),$ (7)
where $\textbf{k}[t],\textbf{v}[t]$ are input-dependent variables called as queries and keys, and $f$ is specific for a certain linear attention. For example, the gated linear attention is:
$\textbf{S}[t]=\textbf{G}[t]\odot\textbf{S}[t-1]+\textbf{v}[t]^{\rm{T}}\textbf{k}[t],$ (8)
where $G[t]$ is the input-dependent gate and $\odot$ is the element-wise multiplication. We modify Eq.(7) as bi-directional, concatenate the forward directional $\textbf{S}_{f}[t]$ and backward direction $\textbf{S}_{b}[t]$, and use a linear layer with weight $\textbf{W}_{o}$ to fuse them to generate the final output $\textbf{O}[t]$:
$\textbf{S}_{f}[t]$ $=f(\textbf{S}_{f}[t-1],\textbf{k}[t],\textbf{v}[t]),$ (9)
$\textbf{S}_{b}[t]$ $=f(\textbf{S}_{b}[t-1],\textbf{k}[L-t-1],\textbf{v}[L-t-1]),$ (10)
$\textbf{S}[t]$ $=[\textbf{S}_{f}[t],\textbf{S}_{b}[t]],$ (11)
$\textbf{O}[t]$ $=\textbf{W}_{o}\textbf{S}[t].$ (12)
The structure of the backbone is shown in Fig.2, which contains NN sequential linear attention blocks. Each block is composed of an average pooling layer, a gated linear attention, and a feedforward network with two linear layers and residual connections [35].
4.2 Event Sampling
The number of events in one sample is extremely large, and the lengths in different samples vary in a huge range, e.g., from 35,267 to 1,594,557 in the DVS Gesture training dataset. Neural networks process inputs in a batched computational fashion, which implies that tensors in one batch should have the identical shape. Thus, we sample $L$ events from each event stream. When $L$ is larger than the number of events in one stream, then it will be padded, and the padding mask will indicate to neural networks to ignore the padded events.
For each event stream in the DVS Gesture dataset, we randomly sample without replacement during training. Randomly sampling with replacement can hardly ensure that all events are sampled. For example, the average number of events in each sample in the DVS Gesture dataset is about 362,239. According to [36], if we randomly sample 4096 events from an event stream with 362,239 events, the expectation of sampling times to cover all events is about 900, indicating that randomly sampling with replacement is inefficient to utilize data. For the ASL-DVS dataset, we empirically find that randomly sampling with replacement works well, even if we have not traversed the whole training set. On the contrary, sampling without replacement requires a huge number of iterations to traverse the training set.
4.3 Training Hyper-parameters
We set all random seeds to 0. We use the AdamW [37] optimizer with the basic learning rate of 0.001, batch size 128, and cosine annealing learning rate scheduler [38]. The actual learning rate is scaled as
$\rm{learning\_rate}=\rm{basic\_learning\_rate}\cdot\rm{batch\_size}\cdot\rm{world\_size}/256,$ (13)
where $\rm{world\_size}=4$ is the number of processes of distributed data parallel training. The specific hyperparameters for different datasets are as follows.
DVS Gesture: We set the embedding dim $D=128$; the convolutional kernel size 3 for the temporal encoding module; the key and value dims $D_{k}=D_{v}=64$, the number of head $n_{head}=4$ for gated linear attentions; $N=8$ linear attention blocks in the backbone with pooling size in each block as $2$, and the hidden feature $D_{ffn}=256$ for the feedforward network. The DVS Gesture dataset contains a small number of samples and is prone to causing overfitting, and we use a series of random data augmentations during training, including removing consecutive events, wrapping consecutive timestamps, resizing, rotation, shear transforming, and translation. We also set the weight decay to 2. The number of training epochs is 8. During training, we randomly sample 8192 events without replacement. During inference, we randomly sample 8192 events from each event stream. Due to the randomness of inference, we report the average test accuracy of repeating 18 times.
ASL-DVS: We set $D=D_{k}=D_{v}=64,n_{head}=2,D_{ffn}=128$. The convolutional kernel size is 3 for the temporal encoding module. We only use $N=1$ linear attention block, and we set the pooling size in the block as $1$ (no pooling). We do not use any data augmentation or weight decay. The number of epochs is 512. During training and inference, we randomly sample 512 events with replacement from each event stream, and the test accuracy is also reported by average over 16 repetitions.
4.4 Analysis of Results
Tab.1 compares the results of our methods and some previous representative methods, where ”Used Training Data (%)” is determined by both training epochs and event data preprocessing methods. For traditional preprocessing methods, such as event-to-frame and randomly sampling without replacement, the ratio is the number of training epochs. For randomly sampling with replacement, each event stream is only sampled once during one epoch. Thus, the ratio is approximately the number of sampled events over the number of events in the training set.
For the DVS Gesture dataset, our method achieves a sub-optimal accuracy compared to Sparse GRU [39] and SNN [40], which may be caused by the miniature scale of the dataset. For the ASL-DLV dataset, our method achieves slightly higher accuracy than previous methods, with a small model size and used training data ratio.
@@ 4.4 Analysis of Results
Dataset Method Preprocessing Accuracy Model Size (MB) Used Training Data (%)
DVS Gesture Sparse GRU [39] To frame 97.80% 4.8 10,000
SNN [40] To frame 97.92% 0.13 19,200
Ours
Randomly sampling
8192 events
96.01% 3.22 800
ASL-DVS GNN+CNN [16] To graph 90.10% 19.46 15,000
GNN + Transformer [41]
To images and
voxel-graph
99.60% 220.3 15,000
Ours
Randomly sampling
512 events
99.80% 0.35 9.7
Table 1: Comparison of methods on DVS Gesture and ASL-DVS datasets.